# Stellar-gas kinematic misalignments in eagle: lifetimes and longevity of misaligned galaxies

### v1

@@ 7 Conclusions
In this work, we have used the eagle cosmological simulations to investigate the longevity of unstable stellar-gas kinematic misalignments in galaxies with mass $>10^{9.5}$ M. We use samples of $\approx 5400$ galaxies for each snapshot to investigate the distribution and variation in misalignment fraction between $10<z<$ 1. We then construct a sample of 3154 misalignment relaxations for which we can reliably estimate the relaxation times. Using this sample, we established the distribution of relaxation timescales for a variety of galaxy properties including morphology, star-forming gas fraction, and environment. Next, we briefly investigated the correlation between the incidence of mergers and the formation of unstable stellar-gas kinematic misalignments. Finally, we compare our results to theoretical predictions and interpret these in the context of driving the observed misalignment distribution at $z\approx 0.1$. Our results can be summarised as follows:
  • For both ETGs and LTGs, the fraction of 3D misalignments ($>30^{\circ}$) increases slightly toward lower redshifts, driven by the build-up of long-lasting counter-rotating ($>150^{\circ}$) systems. By $z\approx 0.1$, we find counter-rotating gas discs present in $\approx 6\%$ of our sample.
  • The majority of unstable misalignments ($[30^{\circ}-150^{\circ}]$) relax on short timescales, with median relaxation times of $t_{\rm{relax}}\approx 0.5$ Gyr $\approx 3.2\bar{t}_{\rm{dyn}}\approx 1.4\bar{t}_{\rm{torque}}$. These indicate the majority of unstable misalignments do not experience significantly enhanced relaxation timescales of $\sim 20\bar{t}_{\rm{dyn}}$ due to continued smooth accretion, as suggested by Davis & Bureau (2016) to explain the shape of the ETG misalignment distribution at $z\sim 0.1$.
@@
  • We find a population ($\approx 20\%$) of long-duration relaxations with $t_{\rm{relax}}>3\bar{t}_{\rm{torque}}$. Galaxies with these long-duration unstable misalignments tend to have lower star-forming gas fractions, above-average stellar mass, and higher ongoing gas inflow rates.
  • We find a positive correlation between halo mass and normalised relaxation timescales, with central galaxies residing in intermediate/high-mass environments ($M_{\rm{200}}>10^{12.5}$ M) showing clear signs of processes preserving unstable misalignments beyond theoretical predictions. We speculate some of these extended lifetimes may be driven by the cooling of hot gas from a misaligned halo.
  • Among galaxies with $M_{*}>10^{10}$ M, mergers coincide with the formation of an unstable misalignment in $\sim 17\%$ of cases in broad agreement with recent observational and simulation work. Major mergers are more effective at forming unstable misalignments and dominate the formation of unstable misalignments ($\approx 12\%$) compared to minor mergers ($\approx 5\%$). Mergers contribute toward the formation of unstable misalignments more in ETGs ($\approx 21\%$) than in LTGs ($\approx 13\%$).
@@
More work needs to be done in this field to fully understand the drivers of these misalignments to discern the main ISM replenishment mechanisms of ETGs. Future observational surveys with e.g. ALMA and JWST alongside upcoming next-generation cosmological hydrodynamical simulations with a multiphase ISM offer an opportunity to explore these unique systems in more detail.
@@ Acknowledgements
The eagle simulations and database are publicly available at: http://icc.dur.ac.uk/Eagle/database.php. Detailed guides to simulation parameters, models, and how to access and query data can be found in The EAGLE team (2017) and McAlpine et al. (2017). eagle "snipshots" are available on request from members of the eagle team. Data and results used in this work can be provided upon request to the author.
@@ Appendix A Variation of latency time on our unstable misalignment relaxation sample
We explore the selection effects of varying the latency time between $0.50<t_{\rm{lat}}<$ 0.5 Gyr, with statistics shown in Table 6. While increasing $t_{\rm{lat}}$ comes at a significant cost to our overall sample size, we find no significant change in the overall shape of our relaxation timescale distribution besides an increase in the relative fraction of $t_{\rm{relax}}<0.3$ Gyr relaxations for $t_{\rm{lat}}=0$ Gyr and subsequently decreased median relaxation time. As expected, this is driven primarily by substantial increase in the relative number of short-duration unstable misalignments rather than a significant decrease in the number of long-duration unstable misalignments. We therefore rule out the possibility of missing any substantial population of long-duration unstable misalignments due to variations in $t_{\rm{lat}}$.
Table 6: Variation of the number of relaxations in our sample (N), median relaxation time ($t_{\rm{relax}}$), and fraction of unstable misalignments with long relaxation times ($t_{\rm{relax}}>1$ Gyr) in our relaxation sample from $10<z<$ 1 with respect to varying latency time ($t_{\rm{lat}}$) between 0 Gyr and 0.5 Gyr.

### v2

@@ 7 Conclusions
In this work, we have used the eagle cosmological simulations to investigate the longevity of stellar-gas kinematic misalignments in galaxies with a stellar mass $10^{9.5}$ M. We use samples of $\approx 5400$ galaxies for each snapshot to investigate the distribution and variation in the fraction of misaligned systems ($\psi_{\mathrm{3D}}>30^{\circ}$) between $0<z<1$. We then construct a sample of 3154 misalignment relaxations for which we can reliably estimate the time taken for an unstable misalignment ($30^{\circ}<\psi_{\mathrm{3D}}<150^{\circ}$) to settle into the kinematically stable regime ($\psi_{\mathrm{3D}}<20^{\circ}$ or $\psi_{\mathrm{3D}}>160^{\circ}$). For each relaxation, we compare the relaxation time ($t_{\mathrm{relax}}$) to the dynamical timescale of the disc ($\bar{t}_{\mathrm{dyn}}$) and to the expected relaxation timescale from theoretical predictions (i.e. the torquing timescale, $\bar{t}_{\mathrm{torque}}$). Using this sample, we established the distribution of relaxation timescales for a variety of galaxy properties including morphology, star-forming gas fraction, and halo mass. Next, we briefly investigated the correlation between the incidence of mergers and the formation of unstable stellar-gas kinematic misalignments. Finally, we compare our results to theoretical predictions and interpret these in the context of driving the observed misalignment distribution at $z\approx 0.1$.
Our results can be summarised as follows:
  1. 1.
    For both ETGs and LTGs, the fraction of 3D misalignments ($30^{\circ}$) increases slightly toward lower redshifts, driven by the build-up of long-lasting counter-rotating ($150^{\circ}$) systems. By $z=0.1$, we find counter-rotating gas discs present in $\approx 6$ per cent of our sample.
  2. 2.
    The majority of unstable misalignments relax on short timescales, with median relaxation times of $t_{\rm{relax}}\approx 0.5$ Gyr $\approx 3.2\bar{t}_{\rm{dyn}}\approx 1.4\bar{t}_{\rm{torque}}$. These indicate the majority of unstable misalignments do not experience significantly enhanced relaxation timescales of $\sim 20\bar{t}_{\rm{dyn}}$ due to continued smooth accretion, as suggested by Davis & Bureau (2016) to explain the shape of the ETG misalignment distribution at $z\sim 0.1$.
  3. 3.
@@
  4. 4.
    We consider long-duration relaxations as those with $t_{\rm{relax}}>3\bar{t}_{\rm{torque}}$ and find that these constitute $\approx 20$ per cent of our sample. Galaxies with these long-duration unstable misalignments tend to have lower star-forming gas fractions, above-average stellar mass, and higher ongoing gas inflow rates.
  5. 5.
    We find a positive correlation between halo mass and normalised relaxation timescales, with central galaxies residing in intermediate/high-mass environments ($M_{\rm{200c}}>10^{12.5}$ M) showing clear signs of processes preserving unstable misalignments beyond theoretical predictions. We speculate some of these extended lifetimes may be driven by the cooling of hot gas from a misaligned halo.
  6. 6.
    Among galaxies with $M_{*}>10^{10}$ M, mergers coincide with the formation of an unstable misalignment in $\approx 10.4$ per cent of cases at $z<0.35$. Major mergers are more effective at forming unstable misalignments and dominate the formation of unstable misalignments ($\approx 7.1$ per cent) compared to minor mergers ($\approx 3.3$ per cent). Mergers contribute toward the formation of unstable misalignments more in ETGs ($\approx 13.9$ per cent) than in LTGs ($\approx 6.4$ per cent). At $0.35<z<1.00$, the incidence of mergers that coincide with misalignment formation increases to $\approx 20.5$ per cent.
@@
More work needs to be done in this field to fully understand the drivers of these misalignments to discern the main ISM replenishment mechanisms of ETGs. Future observational surveys with e.g. ALMA and JWST alongside state-of-the-art cosmological hydrodynamical simulations with a multiphase ISM, such as FIREbox (Feldmann et al., 2023) and the upcoming colibre (Schaye et al., in prep; Chaikin et al. in prep), offer an opportunity to explore these unique systems in more detail.
@@ Acknowledgements
The eagle simulations and database are publicly available at: http://icc.dur.ac.uk/Eagle/database.php. Detailed guides to simulation parameters, models, and how to access and query data can be found in The EAGLE team (2017) and McAlpine et al. (2016). eagle "snipshots" are available on request from members of the eagle team.
@@ Appendix A Variation of latency time on our unstable misalignment relaxation sample
We explore the selection effects of varying the latency time between $0<t_{\rm{lat}}<0.5$ Gyr, with statistics shown in Table 6. While increasing $t_{\rm{lat}}$ comes at a significant cost to our overall sample size, we find no significant change in the overall shape of our relaxation timescale distribution besides an increase in the relative fraction of $t_{\rm{relax}}<0.3$ Gyr relaxations for $t_{\rm{lat}}=0$ Gyr and subsequently decreased median relaxation time. As expected, this is driven primarily by substantial increase in the relative number of short-duration unstable misalignments rather than a significant decrease in the number of long-duration unstable misalignments. We therefore rule out the possibility of missing any substantial population of long-duration unstable misalignments due to variations in $t_{\rm{lat}}$.
Table 6: Variation of the number of relaxations in our sample (N), median relaxation time ($t_{\rm{relax}}$), and fraction of unstable misalignments with long relaxation times ($t_{\rm{relax}}>1$ Gyr) in our relaxation sample from $0<z<1$ with respect to varying latency time ($t_{\rm{lat}}$) between 0 Gyr and 0.5 Gyr.